OpenStudy (rock_mit182):

Which of the following systems of nonlinear equations cannot be solved algebraically ?

OpenStudy (rock_mit182):

A) \[x^2 + 2y^2 =18\] \[xy=4\]

OpenStudy (rock_mit182):

B) \[y=Ln(x) \] \[y=1-x\]

OpenStudy (royalranger):

Second one

OpenStudy (rock_mit182):

c) \[x^2+y^2 = 9\] x^2-y=3

OpenStudy (rock_mit182):

@3mar

OpenStudy (rock_mit182):

please help´

OpenStudy (rock_mit182):

is it B ?

OpenStudy (royalranger):

C) the second one

OpenStudy (rock_mit182):

why C?

OpenStudy (royalranger):

You posted three different questions. Or are they all one big question?

OpenStudy (rock_mit182):

is just one question with three options

OpenStudy (royalranger):

That makes more sense

OpenStudy (rock_mit182):

a b or c ?

OpenStudy (rock_mit182):

@sshayer

OpenStudy (sshayer):

as you can see B can not be solved algebraically

OpenStudy (sshayer):

so B is correct option.

OpenStudy (rock_mit182):

thanks bro I will answer B

OpenStudy (3mar):

I am sorry for late! I was not here!